Can someone help?
What is the slope of the line 15x+3y=90?

Answer by sdmmadam@yahoo.com(530)

15x+3y=90 ----(1)
3y = -15x+90
Dividing through out by 3
y = (-5)x+30 ----(*)
(*) is in the standard slope and y-intercept form:
y = (m)x + b wher m is the slope and b is the y-intercept
Therefore slope of the given line is (-5)
And y-intercept = 30
Note: Given the equation of a straight line
in the general form Ax+By+C=0,
then the slope is given by (-A/B) (if you can remember )
According to this
in the given line (15)x+(3)y-90 =0,
the slope = (-A/B) = -15/3 = -5
Note: Given the line in the general form Ax+By+C=0,
how to convert it into the slope and y-intercept form?
Retain the y-term on the left and take the other terms to the right side.
Divide through out by the coefficient of y.
When the coefficient of y is 1 on the LHS, the coefficient of x on the RHS gives the slope of the line and the free constant now on the RHS
gives the y-intercept.